OpenStudy (anonymous):

Square roots of 2 multiply by square roots of 15

OpenStudy (anonymous):

\[\sqrt{2}* \sqrt{15}\]Is this it?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

you can combine them because there are like terms. What makes them like terms is that both have square roots. You can do this \[\sqrt{2} *\sqrt{15}\]\[\sqrt{2*15}\]

OpenStudy (anonymous):

Can you further simplify that? All you have to do is multiply 15 and 2

OpenStudy (anonymous):

Square roots 30

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yup good job

OpenStudy (anonymous):

If x>2,then x exponent 2 -x -6 is divided by x exponent 2 -4

OpenStudy (anonymous):

If x>2,then x exponent 2 -x -6 is divided by x exponent 2 -4